# Cross-Cultural Expert-Level Art Critique Evaluation with Vision-Language Models

### v2

@@ Abstract
Vision-Language Models (VLMs) excel at visual perception, yet their ability to interpret cultural meaning in art remains under-validated. However, cultural understanding and interpretability are often overlooked when evaluating these models. To overcome this limitation, this paper introduces a tri-tier evaluation framework for cross-cultural art-critique assessment. Tier I provides a series of automated metrics indicating cultural coverage. Tier II leverages theory-informed template-based scoring using a single primary judge across five evaluation dimensions (Coverage, Alignment, Depth, Accuracy, Quality), each rated on a 1–5 scale. Tier III then calibrates the aggregated scores from Tier II via isotonic regression. The proposed evaluation framework is validated with a large-scale experiment covering 15 different VLMs on 294 evaluation art-critique pairs spanning six different cultural traditions. Our findings reveal that (i) automated metrics are unreliable for cultural depth analysis, (ii) Western samples score higher than non-Western samples under our sampling and evaluation template, highlighting potential model biases, and (iii) VLMs exhibit a consistent performance gap, performing well in visual description but underperforming in cultural interpretation. Dataset and code are available at https://github.com/yha9806/VULCA-Framework.
@@
Haorui Yu1, Ramon Ruiz-Dolz2, Xuehang Wen3, Fengrui Zhang4, and Qiufeng Yi5 1DJCAD, University of Dundee, United Kingdom 2Faculty of Science, Engineering and Business, University of Dundee, United Kingdom 3Hebei Academy of Fine Art, China 4Computer Science, Nanjing University, China 5School of Computer Science, University of Birmingham, United Kingdom
@@ 1 Introduction
Vision-Language Models (VLMs) achieve high scores on object-existence probing and VQA-style benchmarks (e.g., POPE Li et al. (2023); VQAv2 Goyal et al. (2017)). Yet these metrics mask a critical limitation: VLMs excel at identifying what is there but fail to interpret what it means. In this paper, we operationalize cultural understanding through five levels: L1 (visual perception), L2 (technical analysis), L3 (cultural symbolism), L4 (historical context), and L5 (philosophical aesthetics). Figure 1 illustrates this limitation, while a VLM correctly identifies “mountains and mist” (L1), it completely misses the Daoist philosophy of emptiness (L5) that gives the painting its cultural significance.
Input: Yuan dynasty landscape painting VLM: “Mountains in mist. Ink wash with varying tones.” Expert: “Sparse brushwork embodies yi (意); pine symbolizes scholarly integrity (L3); composition follows Ni Zan’s dry-brush tradition from Yuan literati school (L4); empty space reflects Daoist wu (无) philosophy (L5)…” Gap: VLM stops at L1–L2; Expert spans L3–L5
Figure 1: VLMs describe visual features (L1–L2) but miss cultural meaning (L3–L5). The expert critique explicitly covers symbolism (L3: pine = integrity), historical context (L4: Ni Zan, Yuan literati), and philosophy (L5: Daoist emptiness).
This limitation has real-world consequences, since VLMs are being widely integrated in many different use cases. Museums deploy VLM-powered guides, art education uses VLMs for critique assistance, and cultural heritage digitization relies on automated annotation. If VLMs mainly focus on superficial aspects of art, and systematically misinterpret non-Western art Cabello et al. (2023); Wu et al. (2025), these applications will perpetuate cultural misunderstanding at scale.
@@
Existing benchmarks fail to capture cultural understanding for two major reasons. First, they assess perception without interpretation, benchmarks like MME Fu et al. (2025) and SEED-Bench-2 Li et al. (2024) test whether VLMs see objects (L1) but not whether they understand symbolism (L3) or philosophy (L5). Second, evaluation methods themselves can sometimes be unreliable. Dual-judge averaging suffers from systematic disagreement, we observe cross-judge Intraclass Correlation Coefficient (ICC) =0.50=-0.50, indicating that judges often operate on incompatible scales. Furthermore, automated metrics cannot substitute for human judgment, with correlations ranging only $r\in[0.27,0.53]$. In culturally-focused generative tasks, evaluation represents a major challenge. If the measuring instrument is unstable across judges and cultures, evaluation rankings become unstable and difficult to reproduce. The central challenge is therefore not only what to measure (e.g., cultural depth with L1-L5) but how to measure it with a controllable and validated instrument. Recent work Yu et al. (2025a, c) examined single cultures, finding significant VLM-expert divergence on Chinese painting critique. But do these failures reflect Chinese art complexity or systematic Western bias? Cross-cultural evaluation is needed to disentangle these factors.
To address these challenges, we introduce a Tri-Tier evaluation built on the L1–L5 cultural schema: Tier I involves automated coverage metrics, Tier II involves template-based single-judge scoring across five evaluation dimensions (i.e., Dimension Covering Rrate, Cultural Semantic Alignment, Critique Depth Score, Factual Accuracy Rate, Linguistic Quality Score), each rated on a 1-5 Likert scale, and Tier III addresses final calibration with human expert references. To avoid confusion, we refer to the evaluation stages with “Tier” and to the cultural understanding depth with “Level” (L1–L5).
This paper addresses three main research questions:
  1. RQ1:
    To what extent can VLMs assist in art critique and cultural understanding across L1-L5?
  2. RQ2:
    How large is the gap between Tier I and Tier II, and can calibration with human references close it?
  3. RQ3:
@@
    Do VLMs exhibit systematic cultural biases in art critique tasks?
Our main contribution is therefore the evaluation framework itself, focusing on the cultural evaluation process rather than a new model architecture or training technique. The specific contribution of this paper is threefold: (1) a Tri-Tier framework combining automated metrics, single-judge scoring, and human reference calibration across 165 culture-specific dimensions (six evaluated traditions); (2) a set of evaluation assets including rubrics, prompts, and diagnostic scripts; and (3) an empirical analysis of 15 different VLMs revealing cultural understanding limitations and biases.
@@ 2.1 VLM Evaluation Benchmarks
VLM evaluation has progressed through three generations. First-generation benchmarks (VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), POPE Li et al. (2023), Winoground Thrush et al. (2022)) assess perceptual grounding at L1. Second-generation benchmarks (MME Fu et al. (2025), SEED-Bench-2 Li et al. (2024), MM-Vet Yu et al. (2024)) introduce multi-task evaluation, but “knowledge” remains encyclopedic rather than cultural.
Third-generation cultural probes reveal gaps: HallusionBench Guan et al. (2024), CulturalBench Chiu et al. (2025), and CulturalVQA Nayak et al. (2024) use multiple-choice QA rather than generative critique; GIMMICK Schneider et al. (2025) documents Western biases but employs recognition tasks rather than interpretation. None operationalize L1–L5 hierarchical understanding or provide frameworks for generative cultural critique.
@@ 2.2 Cultural AI and Bias Studies
Knowledge graph approaches such as WuMKG Wan et al. (2024) for Chinese culture and CulTi Yuan et al. (2026) for general cultural knowledge enable retrieval-based evaluation but do not cover generative evaluation of cultural elements. On the other hand, bias studies document disparities: Cabello et al. (2023) find Western-centric training correlates with non-Western performance degradation, and Wu et al. (2025) quantify demographic representation gaps. However, these studies measure classification accuracy rather than the depth of VLMs’ cultural understanding.
Prior work on VLM art critique reveals specific types of failure. Yu et al. (2025a) evaluated Chinese painting critique using 163 expert commentaries, finding VLM-expert divergence on qiyun shengdong (spirit resonance) interpretation. Yu et al. (2025c) documented “symbolic shortcuts”, i.e., VLMs correctly identify Western Christmas imagery but fail on non-Western festivals like Diwali. These single-culture studies cannot disentangle culture-specific challenges from systematic Western bias, highlighting the need for cross-cultural evaluation following a consistent methodology.
2.3 Computational Aesthetics
Style classification (e.g., WikiArt Saleh and Elgammal (2016) and OmniArt Strezoski and Worring (2018)), and affective computing-based approaches (e.g., ArtEmis Achlioptas et al. (2021)) mostly focus at L1-L2 levels without considering cultural interpretation in-depth. On the other hand, cultural modelling is still in its infancy, Zhang et al. (2024) apply Xie He’s Six Principles but only considering visual and technical aspects (L1-L2), and Jing (2023) theorizes yijing without conducting an evaluation of VLMs on this task. Our framework extends to cultural, historical, and philosophical aspects (i.e., L3-L5) through 165 culture-specific dimensions.
2.4 LLM-as-Judge Paradigm
Finally, when looking at LLM-as-judge approaches such as G-Eval Liu et al. (2023), MT-Bench Zheng et al. (2023), or Prometheus-Vision Lee et al. (2024), they provide scalable evaluation. However, we can observe that cultural critique tasks exhibits three systematic sources of error that dual-judge averaging cannot address: (1) strong judge disagreement (ICC =0.50=-0.50, Claude-Opus vs. GPT-5, $n=150$), (2) incompleteness (specifically, the absence of culture-specific dimensions for concepts such as Chinese wu or Japanese ma), and (3) distribution shift arising from Western-dominant training Tao et al. (2024).
@@ 2.4 LLM-as-Judge Paradigm
Table 1 compares the approach presented in this paper with prior work. We uniquely provide cross-cultural scope spanning 6 different cultural traditions with 165 culture-specific dimensions, hierarchical evaluation considering a deeper set of cultural aspects (i.e., L1-L5) rather than focusing on the surface (i.e.,L1-L2), a validated tri-tier framework with human-focused calibration that measurably improves judge-human alignment and evaluation stability, and systematic bias analysis with a quantified gap between Chinese and Western cultures.
@@
CulturalBench 5 L3 (QA) Auto only
Zhang et al. (2024) Chinese L1–L2 None
Yu et al. (2025a) Chinese L1–L5 Expert
@@ 3 Methodology
We propose a new evaluation methodology to assess cross-cultural art critique generation using a tri-tier framework (see Figure 2) that combines automated metrics, single-judge evaluation, and human-focused calibration.
Evaluation MethodologyInput: VLM Critique ccTier I: Auto Metrics(General Quality Indicators)DCRCSACDSLQSTier II: LLM Judge(Claude Opus 4.5)CovAlignDepthAccQualTier III: CalibrationIsotonic Reg. $g(\cdot)$$S_{\text{II}}^{*}=g(S_{\text{II}})$Cultural Understanding HierarchyL1 VisualL2 TechniqueL3 SymbolismL4 ContextL5 Philosophy  Primary path - - - Suplementary quality indicator
Figure 2: Vulca-Bench Evaluation Framework. Left: Evaluation Methodology where Tier I and Tier II (LLM judge) process the input in parallel. Tier II scores are calibrated considering human-expert references through Tier III producing $S_{\text{II}}^{*}$. Right: L1-L5 cultural understanding hierarchy.
@@ 3.1 Tri-Tier Framework
The evaluation framework processes a VLM-generated critique $c$ for culture $k$ through three tiers. Tier I and Tier II operate in parallel: Tier I computes four automated quality indicators, while Tier II produces template-based scores on five different cultural dimensions from a single primary judge. Then, Tier III calibrates the Tier II aggregate score to human expert ratings via isotonic regression.
@@
The framework receives as an input (i) a bilingual VLM-generated critique asked to cover L1-L5 levels of cultural understanding for a specific artwork image, and (ii) a culture tag. Then produces as an output (a) the calibrated aggregate score ($S_{\text{II}}^{*}$), (b) a template-based assessment covering five different dimensions, and (c) automated metrics-based risk indicators flagging potential types of failure (e.g., low cultural coverage, weak semantic alignment, or elevated template risk).
@@
Tier I evaluation provides information about different aspects of an art critique based on simple heuristics. While they cannot replace a complete expert perspective, they provide a quantitative signal about potential limitations or weaknesses of a given critique. Therefore, we consider the following four automated metrics for a given VLM-generated critique $c$ and a specific culture $k$.
@@
where $v_{c}$ and $v_{k}$ are the TF-IDF vector representations of the critique and the culture-specific vocabulary respectively.
@@
$\text{CDS}_{\text{auto}}(c)=\sum_{\ell=1}^{5}w_{\ell}\cdot\mathbb{1}\text{L}\ell\text{ covered}]$ (3)
where $\mathbb{1}[\cdot]$ is the indicator function (1 if level $\ell$ is detected via keyword matching, 0 otherwise), and $w_{\ell}=\ell/15$ assigns weights L1:1/15, L2:2/15, …, L5:5/15. A critique covering all five levels will achieve CDS=1.0 but one limited to L1–L2 will obtain a score of only 0.2, penalising critiques that focus on visual and technical aspects instead of providing deeper cultural understanding.
@@
$\text{LQS}_{\text{auto}}(c)=\min\left(1,\frac{|c|}{L_{\max}}\right)\cdot\frac{n_{\text{sent}}}{n_{\text{sent}}+\epsilon}$ (4)
where $|c|$ is the character count of critique $c$, $L_{\max}=2000$ is the expected length for an expert critique, $n_{\text{sent}}$ is the number of sentences, and $\epsilon=3$ is a smoothing constant. The first term rewards adequate length and the second term rewards multi-sentence structure. A long, single-sentence response will scores lower, flagging a potential template-style output.
Once calculated, all metrics are rescaled to $[1,5]$ so that they can be compared directly and aggregated within the same scale.
@@
An independent judge (e.g., an LLM) scores each VLM critique following a template addressing five dimensions: Coverage (L1-L5 breadth), Alignment (culture-specificity), Depth (L3-L5 interpretation quality), Accuracy (factual correctness), and Quality (coherence). Each dimension receives a 1-5 rating which is then aggregated as their mean:
@@
We adopt a single-judge design to ensure same-scale outputs. The LLM judge evaluates a critique based on the five-dimension template, then scores are calibrated via Tier III isotonic regression.
Tier III (Calibration).
Tier III calibration aligns the aggregate $S_{\text{II}}$ scores with human judgements via isotonic regression Niculescu-Mizil and Caruana (2005):
$S_{\text{II}}^{*}=\arg\min_{g\in\mathcal{G}}\sum_{i}(g(S_{\text{II}}^{(i)})-S_{\text{h}}^{(i)})^{2}$ (6)
where $\mathcal{G}$ is the set of monotonic functions and $S_{\text{h}}$ is the human score. Isotonic regression corrects systematic judge-human differences on the aggregate score. Dimension-specific scores, however, remain uncalibrated (Tier II) but are also taken into account for analytical purposes. Beyond reducing absolute error, calibration also enables more direct comparability across models and cultures by stabilising rankings under an explicit measurement scale based on human references.

### v3

@@ Abstract
Vision-Language Models (VLMs) excel at visual description yet remain under-validated for cultural interpretation. Existing benchmarks assess perception without interpretation, and common evaluation proxies, such as automated metrics and LLM-judge averaging, are unreliable for culturally sensitive generative tasks. We address this measurement gap with a tri-tier evaluation framework grounded in art-theoretical constructs (§2). The framework operationalises cultural understanding through five levels (L1–L5) and 165 culture-specific dimensions across six traditions: Tier I computes automated quality indicators, Tier II applies rubric-based single-judge scoring, and Tier III calibrates the aggregate score to human expert ratings via sigmoid calibration. Applied to 15 VLMs across 294 evaluation pairs, the validated instrument reveals that (i) automated metrics and judge scoring measure different constructs, establishing single-judge calibration as the more reliable alternative; (ii) cultural understanding degrades from visual description (L1–L2) to cultural interpretation (L3–L5); and (iii) Western art samples consistently receive higher scores than non-Western ones. To our knowledge, this is the first cross-cultural evaluation instrument for generative art critique, providing a reproducible methodology for auditing VLM cultural competence. Framework code is available at https://github.com/yha9806/VULCA-Framework.
@@
Haorui Yu1, Xuehang Wen2, Fengrui Zhang3, and Qiufeng Yi4 1DJCAD, University of Dundee, United Kingdom 2Hebei Academy of Fine Art, China 3Computer Science, Nanjing University, China 4Department of Mechanical Engineering, School of Engineering, University of Birmingham
@@ 1 Introduction
Vision-Language Models (VLMs) achieve high scores on perception benchmarks, yet these metrics mask a critical limitation: VLMs describe what they see but rarely interpret what it means. Art understanding requires moving from visual description (L1–L2) to cultural interpretation, including symbolism, historical context, and philosophical aesthetics (L3–L5, see §2 for theoretical grounding). Figure 1 illustrates this gap: a VLM produces generic praise for a Ming-dynasty farewell painting, while an expert critique identifies specific brushstroke techniques, historical conventions, and aesthetic principles.
Figure 1: VLM vs. Expert on Shen Zhou’s Farewell at Jingjiang. The VLM stops at generic L1–L2 description (score 2.45), while the expert spans L1–L5 with specific technique terms, historical context, and aesthetic interpretation.
@@
Existing evaluation fails to capture this gap for two reasons. First, benchmarks assess perception without interpretation: POPE Li et al. (2023), VQAv2 Goyal et al. (2017), MME Fu et al. (2025), and SEED-Bench Li et al. (2024) test whether VLMs see objects (L1) but not whether they understand symbolism (L3) or philosophy (L5). Second, evaluation methods themselves are unreliable: dual-judge averaging suffers from systematic disagreement (we observe cross-judge ICC(2,1) =0.50=-0.50), and automated metrics correlate only weakly with human judgement. Recent single-culture studies Yu et al. (2025); Yu, Zhao, et al. (2025) find significant VLM-expert divergence on Chinese painting critique, but cannot disentangle culture-specific challenges from systematic Western bias. The central challenge is therefore not only what to measure but how to measure it reliably.
We address this measurement gap with a tri-tier evaluation framework. Tier I computes automated quality indicators that detect whether relevant cultural concepts are mentioned. Tier II applies single-judge rubric scoring that additionally assesses whether cultural properties are correctly attributed and interpreted. Tier III calibrates the aggregate score to human expert ratings via sigmoid calibration. We refer to evaluation stages as “Tiers” and to cultural understanding depth as “Levels” (L1–L5) throughout.
This paper is organised around three research questions:
  1. RQ1.
    Measurement. How should VLM cultural understanding be measured beyond perception-level benchmarks?
  2. RQ2.
    Reliability. How reliable are current evaluation proxies for cultural interpretation?
  3. RQ3.
@@
    Diagnosis. Do VLMs exhibit systematic cultural understanding gaps and biases?
Our contributions, each a verifiable answer to the corresponding question:
  1. C1.
    Construct definition (RQ1): a tri-tier evaluation framework with 165 culture-specific dimensions across six traditions, anchored to expert judgement.
  2. C2.
    Measurement validation (RQ2): empirical evidence that dual-judge averaging and automated proxies are unreliable for cultural evaluation, establishing single-judge sigmoid calibration as the reliable alternative.
  3. C3.
    Diagnostic findings (RQ3): applying the validated instrument to 15 VLMs across six cultural traditions, we find that cultural understanding degrades from L1–L2 to L3–L5 (Depth ($\sigma{=}0.56$) and Alignment ($\sigma{=}0.48$) most discriminative), with 13 of 15 models assigning higher scores to Western art (Cohen’s d=0.74d=-0.74, $p<0.001$).
@@ 2.1 VLM Evaluation Benchmarks
VLM benchmarks have progressed from perceptual grounding (VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), POPE Li et al. (2023)) to multi-task evaluation (MME Fu et al. (2025), SEED-Bench Li et al. (2024)), but assess L1–L2 perception rather than cultural interpretation. Third-generation cultural probes such as CulturalBench Chiu et al. (2025) (text-only QA, 45 regions), CulturalVQA Nayak et al. (2024) (VLMs, 11 countries), and GIMMICK Schneider et al. (2025) (144 countries) test factual cultural knowledge but not generative critique, and none span the full L1–L5 hierarchy.
2.2 Art Interpretation Theory and Computational Approaches
Our framework draws on two complementary traditions in the philosophy of art. Panofsky’s (1939) iconological method distinguishes pre-iconographic description (recognising pure forms and objects), iconographic analysis (identifying conventional themes and symbols), and iconological interpretation (grasping intrinsic cultural and philosophical meaning). We operationalise this ascending hierarchy—from perception through convention to interpretation—as five levels that isolate empirically separable VLM competencies. L1 (visual perception) captures the sensory recognition at the core of Panofsky’s pre-iconographic stratum, while L2 (technical analysis) foregrounds medium and technique identification that practical observation presupposes yet that demands domain training. L3 (cultural symbolism) targets the conventional subject-matter decoding of iconographic analysis, while L4 (historical context) isolates art-historical attribution, which Panofsky distributes across his second and third strata. L5 (philosophical aesthetics) corresponds to iconological interpretation. Goodman’s (1968) theory of symbolic reference motivates the tier distinction: Tier I targets denotational signals and Tier II additionally assesses exemplification and expression.
Computationally, style classification (WikiArt Saleh and Elgammal (2016), OmniArt Strezoski and Worring (2018)) and affective computing (ArtEmis Achlioptas et al. (2021)) address L1–L2; recent VLM approaches target art history comprehension Strafforello et al. (2024) and formal visual analysis Bin et al. (2024) but not cultural interpretation. Zhang et al. (2024) survey computational approaches to Chinese painting through Xie He’s Six Principles (c. 550) but find implementations limited to L1–L2, and Jing (2023) theorises yijing without VLM evaluation. Our framework extends to L3–L5 through culture-specific dimensions (§4).
@@ 2.3 Cultural AI and Bias Studies
Structured knowledge resources such as WuMKG Wan et al. (2024) for Chinese painting and calligraphy and CulTi Yuan et al. (2026) for cultural heritage retrieval enable factual lookup but not generative critique. Bias studies find Western cultural bias in LLM outputs Tao et al. (2024) and fairness disparities across demographic attributes in VLMs Wu et al. (2025), though at survey level rather than depth of understanding. Prior single-culture VLM art critique work includes Yu et al. (2025), who found VLM-expert divergence on qiyun shengdong (spirit resonance), and Yu, Zhao, et al. (2025), who documented symbolic shortcuts on non-Western fire traditions. However, single-culture designs cannot disentangle culture-specific challenges from systematic bias, motivating our cross-cultural approach.
@@
LLM-as-judge approaches (G-Eval Liu et al. (2023), MT-Bench Zheng et al. (2023), Prometheus-Vision Lee et al. (2024)) provide scalable evaluation but face three challenges in cultural critique: strong cross-judge disagreement (§5.2), absent culture-specific vocabularies, and Western-dominant training distribution. Table 1 situates our work: we uniquely combine cross-cultural scope (six traditions, 165 dimensions), full L1–L5 evaluation, and human-calibrated scoring. In sum, prior work probes factual cultural knowledge through closed-form recognition. Our framework instead targets open-ended generative critique and provides measurement validity through human calibration.
@@
CulturalVQA 11 L3 (VQA) Auto only
GIMMICK 144 L3 (recog.) Auto only
Yu et al. (2025) Chinese L1–L5 Expert
@@ 3 Methodology
We propose a tri-tier evaluation methodology for cross-cultural art critique assessment (see Figure 2). Following the theoretical grounding in §2, the design reflects the key distinction between surface-level lexical signals and deeper cultural understanding: automated metrics capture keyword presence, while judge-based scoring assesses whether cultural concepts are correctly attributed and interpreted. The three tiers combine automated metrics, single-judge evaluation, and human-focused calibration to address this distinction systematically.
[…]
Figure 2: Vulca-Bench Evaluation Framework. The L1–L5 cultural understanding hierarchy (centre) is the construct being measured. Coloured span bars show which levels each metric assesses: DCR, CSA, Coverage, and Accuracy span L1–L5 (full bars), while CDS, Depth, and Alignment target L3–L5 (darker bars). LQS and Quality are level-agnostic (dotted). The dashed line marks the surface/deep boundary. The Tier II aggregate is calibrated to human ratings via Tier III, producing $S_{\text{II}}^{*}$.
@@ 3.1 Tri-Tier Framework
We distinguish Levels (L1–L5) as the construct under measurement from Tiers (I–III) as the measurement mechanisms applied to that construct (see Figure 2 and Table 2). The evaluation framework processes a VLM-generated critique $c$ for culture $k$ through three tiers. Tier I and Tier II operate in parallel: Tier I computes four automated quality indicators, while Tier II produces rubric-based scores on five different cultural dimensions from a single primary judge. Then, Tier III calibrates the Tier II aggregate score to human expert ratings via sigmoid calibration.
@@
The framework receives as input (i) a bilingual (Chinese–English) VLM-generated critique asked to cover L1–L5 levels of cultural understanding for a specific artwork image, (ii) a culture tag, and (iii) an expert reference critique for the same artwork. Bilingual output is required because Chinese preserves culture-specific terminology essential for keyword-based Tier I metrics (e.g., qiyun 气韵), while English ensures cross-cultural accessibility. Tier I metrics use only (i) and (ii). Tier II operates in reference-guided mode, where the judge compares the VLM critique against the expert reference to assess cultural alignment and factual accuracy. The framework produces as output (a) the calibrated aggregate score ($S_{\text{II}}^{*}$), (b) a rubric-based assessment covering five dimensions, and (c) automated risk indicators flagging potential failure types (e.g., low cultural coverage, weak semantic alignment, or elevated template risk).
@@
Tier I captures surface-level lexical signals: it detects whether a critique mentions relevant cultural concepts through keyword and vocabulary matching, without assessing whether those concepts are correctly attributed or interpreted. While these heuristics cannot replace expert perspective, they provide a quantitative signal about potential limitations. We define four automated metrics for a VLM-generated critique $c$ and culture $k$.
@@
where $\mathbf{v}_{c}$ and $\mathbf{v}_{k}$ are the TF-IDF vector representations of the critique and the culture-specific vocabulary respectively.
@@
$\text{CDS}_{\text{auto}}(c,k)=\sum_{\ell=1}^{5}w_{\ell}\cdot\mathbb{1}[\text{L}\ell\text{ covered}]$ (3)
where $\mathbb{1}[\cdot]$ is the indicator function (1 if level $\ell$ is detected via keyword matching, 0 otherwise), and $w_{\ell}=\ell/15$ assigns weights L1:1/15, L2:2/15, …, L5:5/15.
@@
$\text{LQS}_{\text{auto}}(c,k)=\min\left(1,\frac{|c|}{L_{\max}}\right)\cdot\frac{n_{\text{sent}}}{n_{\text{sent}}+\epsilon}$ (4)
where $|c|$ is the character count, $L_{\max}=2000$ the expected expert-critique length, $n_{\text{sent}}$ the sentence count, and $\epsilon=3$ a smoothing constant.
Each metric is rescaled to $[1,5]$ via $\hat{m}=1+4\cdot\min(1,\max(0,m))$ and the Tier I aggregate is their mean: $S_{\text{I}}=\frac{1}{4}\sum_{m\in\mathcal{M}_{\text{I}}}\hat{m}$. Tier I signals serve as diagnostic risk indicators rather than primary ranking metrics.
@@
Unlike Tier I, Tier II assesses deeper cultural understanding: the judge evaluates not only whether concepts are mentioned but whether cultural properties are correctly attributed and whether deeper aesthetic meanings are appropriately interpreted. In reference-guided mode (defined above), the judge scores five dimensions: Coverage (L1–L5 breadth), Alignment (culture-specificity), Depth (L3–L5 interpretation quality), Accuracy (factual correctness), and Quality (coherence). Each dimension receives a 1–5 rating, and the aggregate score is their mean:
@@
where $d_{i}$ is the score for dimension $i$.
We adopt a single-judge design to ensure same-scale outputs. The judge evaluates each VLM critique against the paired expert reference using the five-dimension rubric, and the aggregate score is then calibrated via Tier III sigmoid calibration. The framework also supports a reference-free mode (rubric-only, no expert anchor) for deployment settings where references are unavailable. All results in this paper use reference-guided evaluation. Crucially, the expert critique serves as an alignment anchor rather than a unique gold answer, since L3–L5 dimensions admit legitimate interpretive plurality.
Table 2 summarises how each evaluation component maps onto the L1–L5 hierarchy, making the construct–mechanism relationship explicit.
Component L1–L2 L3–L5
Tier I (Automated)
   DCR
   CSA
   CDS
   LQS
Tier II (Judge)
   Coverage
   Accuracy
   Depth
   Alignment
   Quality
Tier III
   Calibration aggregate only
Table 2: Tier–Level mapping. Checkmarks indicate which cultural understanding levels each component explicitly assesses. L1–L2 covers visual perception and technical analysis; L3–L5 covers cultural symbolism, historical context, and philosophical aesthetics.
Agreement Reporting (ICC).
For agreement analyses, we use ICC(2,1) (two-way random effects, absolute agreement, single measurement) following reporting guidance in Koo & Li (2016) and McGraw & Wong (1996) Koo and Li (2016); McGraw and Wong (1996). We report 95% confidence intervals for Tier I–Tier II ICC estimates (Appendix Table 8). We use ICC rather than Pearson/Spearman correlation because our target is interchangeability under a shared scale (absolute agreement), while correlation measures association and can remain high under additive or scaling mismatches.
Tier III: Human Calibration.
Tier III grounds the evaluation in human expert judgement, addressing the concern that even judge-based scoring may exhibit systematic biases. Calibration aligns $S_{\text{II}}$ (Eq. 5) with human ratings via sigmoid calibration Niculescu-Mizil and Caruana (2005):
$S_{\text{II}}^{*}=1+4\,\sigma(a\cdot S_{\text{II}}+b)$ (6)
where $\sigma(\cdot)$ is the logistic function and $a,b$ are fitted by minimising MSE on the training split of human-scored pairs ($n=295$), with $S_{\text{h}}$ denoting the human score. The sigmoid maps the unbounded judge aggregate onto $[1,5]$ while preserving monotonicity by construction. Dimension-specific scores remain uncalibrated (Tier II) but are reported for analytical purposes. Beyond reducing absolute error, calibration enables comparability across models and cultures by anchoring rankings to an explicit measurement scale grounded in human expertise, a requirement for cross-cultural claims where systematic bias is itself under investigation.